OpenStudy (anonymous):

PLEASE HELP!!! What is the simplified form of 5/3y+2 - 4/9y^2-4

OpenStudy (anonymous):

a) 15y-6/(3y-2)(3y+2) b) 15y-6/(3y+2)(3y+2) c) 15y-14/(3y+2)(3y-2) d) 15y-14/(3y+2)(3y+2)

OpenStudy (zzr0ck3r):

9y^2-2 = (3y-2)(3y+2) so 5/3y+2 - 4/9y^2-4 = 5/(3y+2) - 4/((3y+2)(3y-2)) so we multiply the left by (3y-2) on top and bottom so that we can have a common denominator (5(3y-2)-4)/((3y-2)(3y+2))

OpenStudy (zzr0ck3r):

can you simplify?

OpenStudy (zzr0ck3r):

just simplify the top 5(3y-2)-4

OpenStudy (zzr0ck3r):

@danny562 ?

OpenStudy (zzr0ck3r):

you should participate :)

jhonyy9 (jhonyy9):

@zzr0ck3r ATTENTION !!! in your first answer there need being -4 and not -2 !!! ok ? correct it please !!!

OpenStudy (zzr0ck3r):

right 9y^2-4 = (3y)^2-2^2 = (3y-2)(3y+2)

jhonyy9 (jhonyy9):

@danny562 so do you understand it ,can you solve it now ?

jhonyy9 (jhonyy9):

so just continue it with this 5(3y-2)-4 = ?

jhonyy9 (jhonyy9):

so = 15y-10 -4 = 15y -14 and for denominator will be (3y-2)(3y+2) so if you check these choises you see what will be right ... ok ? wait your answer ---

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

???

jhonyy9 (jhonyy9):

are you here now ?

OpenStudy (anonymous):

5/3y+2-4/9y^2-4 or,5/3y+2-4/3y^2-2^2 or,5/3y+2-4/(3y+2)(3y-2) or,5(3y-2)-4/(3y+2)(3y-2) or,15y-10-4/(3y+2)(3y-2) or15y-14/(3y+2)(3y-2)

jhonyy9 (jhonyy9):

@Spynextdoor ATTENTION !!! in the second line the denominator of second fraction so not is right you have missed there an parentheses not is just y squared the 3 is squared too ok ?